### v4

# Fairshare Data Pricing via Data Valuation for Large Language Models

Abstract
Training data is the backbone of large language models (LLMs), yet today’s data markets often operate under exploitative pricing – sourcing data from marginalized groups with little pay or recognition. This paper introduces a theoretical framework for LLM data markets, modeling the strategic interactions between buyers (LLM builders) and sellers (human annotators). We begin with theoretical and empirical analysis showing how exploitative pricing drives high-quality sellers out of the market, degrading data quality and long-term model performance. Then we introduce fairshare, a pricing mechanism grounded in data valuation that quantifies each data’s contribution. It aligns incentives by sustaining seller participation and optimizing utility for both buyers and sellers. Theoretically, we show that fairshare yields mutually optimal outcomes: maximizing long-term buyer utility and seller profit while sustaining market participation. Empirically when training open-source LLMs on complex NLP tasks, including math problems, medical diagnosis, and physical reasoning, fairshare boosts seller earnings and ensures a stable supply of high-quality data, while improving buyers’ performance-per-dollar and long-term welfare. Our findings offer a concrete path toward fair, transparent, and economically sustainable data markets for LLM.
1 Introduction
High-quality training data is foundational to building effective and reliable large language models (LLMs). As LLMs take on increasingly complex tasks today – such as coding[chen2021humaneval], reasoning [wei2022chain], and AI4Science [taylor2022galactica] – they rely heavily on carefully curated, human-annotated data. This growing demand has triggered a "generative data gold rush", with major tech companies racing to acquire training data, fueling the rise of a nascent AI data market [reuters2024bigtech]. In this market, AI firms create networks of short-term contract workers to generate data labels, resembling an Uber-like gig economy for data [reuters2024bigtech].
However, the current AI data market operates with limited oversight and is widely criticized for a lack of transparency and fairness in pricing [paul2024bigtech, zhang2024surveydatamarkets, metcalf2021algorithmic, akerlof1978market]. Data prices are largely low and fail to reflect the quality or effort involved, threatening the sustainability and quality of data supply [mason2009financialcrowds, hara2018datadrivenamt, cbsnews2024]. In particular, for data sellers, such as human annotators or content creators, the prevailing market routinely undervalues their labor, offering compensation that neglects the skill, effort, and downstream value of their contributions. These harms are especially concentrated in low-wage labor markets, where annotators often face overwork, underpayment, and exclusion from decision-making [time2023kenya]. This reflects a broader ethical concern known as “AI parachuting”, where developers extract data from marginalized communities [jo2020archives, birhane2021impossibility, qz_africa_parachute], tieing to wider debates on epistemic injustice and data colonialism [paullada2021datadiscontents, sambasivan2021datawork, gray2019ghost]. As one civil rights advocate noted on 60 Minutes, “They don’t pay well… they could pay whatever, and have whatever working conditions.” [^1] CBS, 60 Minutes
Motivated by these issues, we present a fair pricing framework for the LLM training data market to promote equitable and sustainable generative AI ecosystems. Economic theory suggests that prices should reflect the value delivered to the buyer – signaling quality and alignment [spence1978job]. Guided by this, our framework introduces fairshare pricing based on established data valuation techniques for LLMs [pruthi2020estimating, kwon2023datainf, xia2024less, yu2024mates], which quantify each dataset’s contribution to model performance. Buyers and sellers both have access to data valuation scores, which guide decision-making on both sides: buyers use them to select datasets under budget constraints to maximize utility (i.e., a standard measure of welfare or satisfaction [mas1995microeconomic, von2007theory]), while sellers use them to set prices based on the anticipated demand from buyers. This shared access – enabled by our assumption of information transparency – supports procedural fairness [konovsky2000understanding], improves seller participation, and enhances overall data quality.
Our theoretical and empirical findings show that fairshare pricing offers clear advantages compared to existing methods. First, we show that existing exploitative pricing leads to a lose-lose outcome for the data market. For data buyers, underpaying data sellers might cut costs in the short term – but it comes at a cost of drive sellers away, shrinking the supply of high-quality training data. This weakens the data pipeline and limits model improvement, even as investments grow. In contrast, we show theoretically that fairshare pricing leads to a win-win outcome: sellers maximize profit while remaining engaged, and buyers secure long-term utility by maintaining access to high-quality data.
Second, we empirically validate our approach through simulations of buyer-seller interactions in data markets. We focus on training open-source LLMs on complex NLP tasks, including math problems [amini2019mathqa], medical diagnosis [jin2020medqa], and physical reasoning [bisk2020piqa]. Analyzing both pricing and valuation outcomes, we find that under fairshare pricing, buyers achieve higher model performance per dollar spent, making it particularly beneficial for those with limited budgets. In addition, our simulations of long-term market dynamics demonstrate that fairshare pricing encourages sustained seller participation, resulting in a stable and sufficient supply of training data over time compared to exploitative pricing. These findings show that our framework’s data-valuation-based pricing not only improves short-term training efficiency, but also ensures the long-term viability of the data market.
Finally, to evaluate the robustness and method-agnostic applicability of our framework, we conduct an ablation study using a diverse set of data valuation methods (including BM25 [trotman2014improvements], $\text{Infl}_{\text{IP}}$ [xia2024less], and Datainf[kwon2023datainf]) – selected for their scalability and efficiency in LLM tasks. Across all variants, our fairshare pricing framework consistently delivers mutually beneficial results for both buyers and sellers in the LLM data market, confirming that its performance is not tied to any specific data valuation technique. This analysis underscores a key strength of our approach: buyers and sellers can flexibly choose valuation methods tailored to their downstream needs without compromising the incentive-aligned structure of the market, and demonstrates the broad applicability of our solution.
To summarize our contributions are as follows:
  1. 1.
    We present a novel theoretical framework to model the LLM training data market utilizing data valuation, economics of market, and game theory.
  2. 2.
    We propose a fairshare pricing mechanism that captures the strategic interplay between buyers and sellers. Both theory and empirical analyses show that exploitative pricing results in lose-lose outcomes, while fairshare promotes long-term stability and mutual benefits.
  3. 3.
    Our proposed pricing framework is highly generalizable across LLM models and tasks, and robust to the data valuation methods and utility measures used by market participants.
Our results demonstrate the versatility of our approach in data markets by aligning incentives, sustaining data quality, and maximizing long-term value. This highlights our fairshare mechanism as a guide for designing sustainable data procurement practices in public and private sectors.
2 Related Work
Data Market Design and Exploitative Pricing. Recent studies on ML data markets focus on platform-based pricing and allocation mechanisms such as auctions and personalized pricing [agarwal2019marketplace, chen2019towards, chen2022selling, yang2022selling, zhang2024survey]. These frameworks typically aim to maximize profit or efficiency, assuming myopic buyers and fixed seller participation. They also consider information transparency ensures that sellers have access to the same data valuation scores used by buyers [ghorbani2020distributional, chen2022selling]. The study was mainly conducted on classic machine learning models rather than large language models.
Concurrently, research has highlighted ethical concerns in LLM data acquisition, particularly around exploitative pricing. Annotation work – central to LLM training – is often outsourced to low-wage workers with minimal labor protections [mason2009financialcrowds, hara2018datadrivenamt, toxtli2021quantifying, remotasks_wapo_2023], who are frequently exposed to harmful content [paul2024bigtech, time2023kenya, weidinger2021ethicalsocialrisksharm]. Several studies document low pay, precarious labor, and lack of protections in global data supply chains [paullada2021datadiscontents, sambasivan2021datawork, gray2019ghost]. Although these studies highlight systemic exploitation in data labor, their findings have yet to be integrated into formal models of LLM data pricing. Economic theory suggests that misaligned compensation undermines efficiency, discourages participation, and lowers data quality. In LLM contexts, this reduces data availability and degrades model performance [techcrunch2024, seetharaman2024gpt5, javed2025aws], underscoring the need for value-aligned pricing mechanisms.
Game-Theoretic Models of Pricing and Participation. In many domains, pricing problems are modeled thorugh game theory. Particularly, those involving decision-making are frequently modeled using Stackelberg games and bilevel optimization [brander1985export]. Although these frameworks have not yet been applied to LLM data market, they have been extensively used in contexts such as resource allocations [bard2013practical], energy market [von2010market, pei2020survey], supply chains [esmaeili2009game], distributed systems [maharjan2013dependable]. In these settings, agents typically optimize revenue, efficiency, or utility while anticipating the strategic responses of others. In dynamic settings, these models extend to repeated interactions, where players balance immediate gains against long-term benefits [stokey1989recursive, brander1985export]. In repeated interactions, fairness perceptions are critical: studies show that perceived pricing unfairness can erode trust and reduce participation [folger1998organizational, folger2001fairness, colquitt2013organizational, adamovic2023organizational].
Data Valuation Methods: Data valuation estimates the contribution of training examples to model performance [hammoudeh2024training]. Examples include influence-function-based methods [koh2017understanding], which estimate data utility via model gradient computations. Variations such of this method enhance efficiency by approximating/bypassing the inverse-Hessian [pruthi2020tracin], or utilizing lower dimension model gradients (e.g., Datainf [kwon2023datainf]), which balance efficiency and accuracy, and make them suitable for the LLM realm. Beyond the influence of loss functions, previous research [wang2022understanding] has examined data valuation with respect to fractional utility functions.
Data valuation approaches are effective for LLM training data selection [xia2024less, yu2024mates, jiao2025datelm] and other applications including toxicity detection [isonuma2024unlearning], memorization analysis [kounavis2023influence], training optimization [van2023hint], and and active sampling [pang2024fairness]. Shapley-based method is another direction [shapley1953value, jia2019towards, ghorbani2019data, ghorbani2020distributional], such as CS-Shaply[schoch2022cs], In-run Shapley[wang2024data], DU-Shapley[garrido2024shapley], etc. Simpler methods like BM25 [trotman2014improvements] offer a model-agnostic baseline based on lexical similarity[akyurek-etal-2022-towards, wu-etal-2024-enhancing-training]. While exact data value estimation is computationally costly at LLM scale, recent studies show that influence-based approximations correlate meaningfully with actual performance outcomes [jiao2024context, wang2024greats]. For example, Jiao et al. [jiao2024context] shows that influence-based methods have high correlation between their estimated data valuation scores and the oracle value, which is obtained through model re-training.
3 A Theoretical Framework of LLM Data Market
This section formalizes the LLM data market. We model buyer–seller interactions using economic utility theory and game theory, show that exploitative pricing leads to long-term market failure, and introduce the fairshare pricing strategy.
3.1 Data Market Definition
We formalize the LLM data market by modeling the decision-making processes of sellers and buyers in a non-cooperative Stackelberg game [zhang2024survey, von2010market, maharjan2013dependable]. In this setup, sellers first set prices by anticipating buyer demand. Buyers then respond by selecting datasets to maximize their utility. [^2] Utility is defined as a general economic concept that measures an individual’s welfare, benefit, or satisfaction [mas1995microeconomic, von2007theory], for example, monetary value, model fairness, or other domain-specific benefits. Each player makes decisions based solely on the current state of the market, without considering future updates or interactions.
Notably, our choice of a non-cooperative Stackelberg game fits the LLM data market’s reality: many independent, self-interested transactions where binding grand coalitions are infeasible, such as gig-work platforms [gray2019ghost]. In contrast, cooperative bargaining targets few-party, contractible surplus splits, which is a different regime and question. In this non-cooperative game, we model fairness behaviorally via seller participation/exit that keeps the model tractable, and the leader–follower timing mirrors real price setting and demand response.
Figure 1: Illustration of the LLM data market, showing buyers (LLM builders) and sellers (annotators) interacting via a pricing mechanism based on game theory and data valuation.
Following prior work [tadelis2016reputation], we assume information transparency: all participants observe each dataset’s value. Practically, this can be estimated using data valuation methods that quantify its contribution to LLM performance. These valuation scores – potentially generated by the platform or a trusted third party – serve as common signals informing both pricing and purchasing decisions.
Next, we define key factors determining each player’s decision-making. In practice, this can be estimated using data valuation methods that quantify its contribution to LLM performance. These scores – potentially generated by the platform or a trusted third party – serve as common signals guiding both pricing and purchasing decisions.
In the rest of this section, we formally define key factors determining each player’s decision-making.
Data Market Setup. We begin by defining the players and the notion of utility. The market comprises a set of buyers $\{B_{k}\}_{k=1}^{M}$, each with an LLM $\mathcal{M}_{k}$, and a set of sellers $\{S_{j}\}_{j=1}^{N}$, each with a dataset $D_{j}$. Each buyer gains a certain level of economic utility from acquiring a new dataset based on a data value, which is captured by a data valuation function, $v_{k}:D\rightarrow\mathbb{R}_{\geq 0}$, measuring the marginal contribution of each dataset $D$ to the performance of buyer $B_{k}$’s LLM $\mathcal{M}_{k}$.
The value $v_{k}$ can be estimated using a data valuation method (e.g., Influence Function [koh2017understanding], DataInf [kwon2023datainf], etc.). Once the data valuation score $v_{k}$ is estimated, the corresponding utility gain $u_{k}$ can be derived as a function of $v_{k}$. We use a task-specific mapping between data value $v_{k}$ and utility $u_{k}:v_{k}\rightarrow\mathbb{R}_{\geq 0}$. Appendix˜C lists common mappings between $u_{k}$ and $v_{k}$ in downstream tasks.
Buyer’s Decision-Making. Each buyer $B_{k}$ selects a subset of datasets to acquire, maximizing the net utility, defined as the utility gain minus total acquisition cost.
$B_{k}$’s decision, represented by a binary vector $\boldsymbol{x}$ (where each entry is 1 if the corresponding dataset is selected, and 0 otherwise), depends on three components: (i) the current dataset prices, (ii) buyer’s budget $b_{k}$, and (iii) the utility gain $u_{k}(\boldsymbol{x})$ from acquiring the selected datasets.
This utility gain reflects the downstream value, resulting from performance improvements in the buyer’s model after training on the dataset bundle $\boldsymbol{x}$. [^3] Datasets may exhibit dependencies. The set utility $u_{k}(\boldsymbol{x})$ is not necessarily additive over individuals. The net utility is defined as:
$g_{k,N}\left(\boldsymbol{x}\right):=u_{k}\left(\boldsymbol{x}\right)-\boldsymbol{x}^{T}\boldsymbol{p},$ (1)
where $\boldsymbol{p}:=[p_{1},\dots,p_{N}]$ is the price vector for current avaliable datasets.
Finally, $B_{k}$’s purchasing problem is formulated as selecting an optimal collection of datasets to maximize its net utility:
$\tilde{\boldsymbol{x}}^{k,N}:=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}_{k,N}}\;g_{k,N}(\boldsymbol{x}),\quad\text{s.t.}\quad\mathcal{X}_{k,N}:=\{\boldsymbol{x}\mid g_{k,N}(\boldsymbol{x})\geq 0,\boldsymbol{x}^{T}\boldsymbol{p}\leq b_{k}\},$ (2)
where $\tilde{\boldsymbol{x}}^{k,N}$ is the optimal solution, and $\mathcal{X}_{k,N}$ includes all feasible solutions, $\boldsymbol{x}^{T}\boldsymbol{p}\leq b_{k}$ ensures that the total acquisition cost is within the budget $b_{k}$, and $g_{k,N}(\boldsymbol{x})\geq 0$ ensures a non-negative net utility.
Seller’s Decision-Making. When the seller $S_{j}$ offers dataset $D_{j}$, it sets a price to maximize its net profit, defined as (i) the anticipated sales from all buyers, minus (ii) a fixed cost to create the dataset. Anticipated sales from buyers are estimated by solving the previous buyer’s purchasing problem (Equation˜2) known in the full transparent market. Formally, the seller’s net profit function is:
$r(p_{j}):=\sum_{k=1}^{M}\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})p_{j}-c_{j},$ (3)
where $\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})$ is $j$-th entry of buyer $B_{k}$’s decision vector, indicating if $B_{k}$ purchases $D_{j}$ ($j$-th entry is $1$) or not ($j$-th entry is 0), given $\boldsymbol{p}$ the prices of all datasets in the market; $c_{j}$ denotes the cost. The seller $S_{j}$ solves the following pricing problem:
$p^{*}_{j}:=\operatorname*{arg\,max}_{{p}_{j}\in\mathcal{P}_{j,M}}\;r(p_{j}),\quad\text{s.t.}\quad\mathcal{P}_{j,M}:=\{{p}_{j}\in\mathbb{R}_{+}\mid r(p_{j})\geq 0\},$ (4)
where $p^{*}_{j}$ is the optimal price and $r(p_{j})\geq 0$ ensures that seller $S_{j}$’s net profit must be non-negative. It is noted that we assume that selling data for annotator is profitable, i.e., $p^{*}_{j}\geq c_{j}$.
In the above formulation, buyer purchase decisions depend on prices and data value; seller pricing decisions depend on anticipated sales from buyers. Together, these decisions define the equilibrium dynamics of a transparent and incentive-aligned LLM data market. This formulation can be extended to a royalty-based scheme as presented in Appendix˜B.
3.2 Exploitative Pricing: A Lose-Lose Outcome
We now present a theoretical analysis showing that exploitative pricing in data markets is ultimately unsustainable and detrimental to all participants.
Theoretical Setup. We analyze a simplified dynamic setting involving a single buyer and a single seller over an infinite time horizon. At each time step $t$, both players decide whether to transact based on the proposed price. Importantly, to align with real-world market behavior, the seller’s future participation depends on the history of past transaction prices.
We first introduce some assumptions on the seller behavior.
Assumption 1 (Declining Participation Probability).
When offered a price $p_{t}$ below the seller’s ideal price $p^{*}_{t}$, the probability of seller’s continued participation declines. This decline is captured by a strictly increasing function
$\pi:(p_{t},p^{*}_{t})\rightarrow[0,1],\text{ satisfying }\pi(0,p^{*}_{t})=0\text{ and }\pi(p^{*}_{t},p^{*}_{t})=1.$ (5)
In addition, the probability that the seller $S$ remains active at time $T$ is $P_{T}:=\prod_{t=0}^{T-1}\pi(p_{t},p^{*}_{t})$.
Assumption˜1 models how the seller respond to exploitative pricing over time. It states that when offered a price below its ideal level, the seller become less likely to stay in the market, consistent with prior studies [gale1994bottom, backus2020sequential]. For any exploitative pricing, the participation probability $P_{T}$ declines multiplicatively, leading to a sustained and irreversible drop in engagement.
To capture the degree of participation decline, we assume sellers respond sensitively to underpayment:
Assumption 1.1 (Sensitivity of Participation Is Lower-Bounded).
The Lipschitz continuity of $\pi$ over exploitative pricing is lower-bounded:
$\left|\pi(p_{t,1},p^{*}_{t})-\pi(p_{t,2},p^{*}_{t})\right|\geq L\left|p_{t,1}-p_{t,2}\right|,$ (6)
for some constant $L>0$ and all exploitative pricing $p_{t,1},p_{t,2}\in[0,p^{*}_{t})$ for all $t$.
We also introduce the following assumptions on buyers:
Assumption 2 (Discount Factor Is Lower-Bounded).
The buyer evaluates long-term utility using a discount factor $\delta$, satisfying:
$\delta\geq\frac{1}{1+L\min_{t\in[0,\infty)}\mathbb{E}[u_{t}-p_{t}^{*}]},\forall t.$ (7)
A higher discount factor $\delta$ indicates greater emphasis on future utility. Assumption˜2 places a lower bound on $\delta$, ensuring that the buyer sufficiently values future gains when making decisions.
Assumption˜2 assumes the buyer values both immediate and future utility gains from acquiring training data. Major LLM developers invest in large-scale data acquisition and model training in expectation of long-term gains in performance, deployment value, and commercial returns [chow2024aws, dotan2024gpt5, li2024time100ai, mayer2025superagency].
(a) Buyer’s utility.
(b) Seller’s participation.
(c) Seller $S_{1}$’s profit.
(d) Seller $S_{2}$’s profit.
Figure 2: Buyer’s cumulative utility and seller participation under ideal and exploitative pricing (2(a),2(b)); Profits of sellers $S_{1}$ and $S_{2}$ with buyer $B_{1}$’s and $B_{2}$’s MWP over price (2(c),2(d)).
Buyer’s Objective. With the previous setup, the buyer aims to maximize expected cumulative utility over an infinite horizon. Let $G(u_{t},b_{t})$ denote this value function, representing the buyer’s expected total cumulative utility at time $t$, conditioned on current utility $u_{t}$ and budget $b_{t}$. Then this objective satisfies the following Bellman equation[bellman1957dynamic]:
$G(u_{t},b_{t})=\max_{p_{t}\in[0,\infty)}\left[\mathbb{E}\left[u_{t}-p_{t}\right]+\delta\mathbb{E}\left[\pi(p_{t},p^{*}_{t})G(u_{t+1},b_{t+1})\mid u_{t},b_{t}\right]\right].$ (8)
This Bellman equation captures the buyer’s central trade-off: offering an exploitative price $p_{t}$ increases immediate surplus $\mathbb{E}[u_{t}-p_{t}]$, but reduces future seller participation via a lower $\pi(p_{t},p^{*}_{t})$; conversely, setting a fairer price decreases short-term gain but sustains future transactions by increasing $\pi(p_{t},p^{*}_{t})$. With this insight, we then obtain the following result:
Lemma 1 (Inevitable Failure of Exploitative Pricing).
With Assumptions˜1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
Lemma˜1 reveals a fundamental limitation: any exploitative pricing strategy is only optimal for a finite time. Over time, it becomes strictly suboptimal, due to the declining seller participation. Thus, no exploitative strategy can maximize cumulative utility in the long run. (See Appendix˜D for the proof, shows that exploitative pricing yields a suboptimal value function.)
Exploitative Pricing Leads to Lose-lose: While exploitative pricing clearly reduces seller welfare, our results show it also harms long-term cumulative utility for buyers. Although buyers may benefit initially from lower costs, reduced seller participation quickly leads to a decline in data quality and availability. Over time, even well-resourced buyers face data shortages. This sets off a self-defeating cycle where short-term savings come at the expense of long-term model performance.
In addition to our theoretical findings, we run a simplified dynamic market simulation with one buyer and one seller, making sequential decisions. The utility of the dataset and buyer’s budget varies randomly over time, and seller participation follows $\pi(p_{t},p_{t}^{*})=p_{t}/p_{t}^{*}$. Figures˜2(a) and 2(b) provides consistent evidence that exploitative pricing causes rapid seller exit and an immediate shortage of training data. This dynamic closely resembles the classic “market for lemons” problem, where underpricing drives out high-quality supply, ultimately leading to market collapse [akerlof1978market].
3.3 Fairshare Pricing: A Win-Win for Sellers and Buyers
The rest of this section presents the fairshare pricing in our LLM data market framework and show it yields a stable, mutually optimal outcome for both buyers and sellers.
Seller-Side: Pricing Based on Buyers’ Maximum Willingness to Pay. When the seller $S_{j}$ prices its dataset $D_{j}$, it evaluates how each buyer’s decision changes depending on whether $D_{j}$ is available at a given price. For each buyer $B_{k}$, the seller compares two scenarios: (1) the buyer’s optimal dataset selection when $D_{j}$ is not part of the market, and (2) the buyer’s new optimal selection if $D_{j}$ is included at price $p_{j}$. The buyer will purchase $D_{j}$ only if the dataset $D_{j}$ increases its net utility and remains within budget. Formally, let $\tilde{\boldsymbol{x}}^{k,N-1}$ be buyer $B_{k}$’s optimal decision without the presence of $D_{j}$. For any feasible decision $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$, we define:
  1. 1.
    Marginal utility from adding $D_{j}$: $\Delta u_{k}(\boldsymbol{x}):=g_{k,N}(\boldsymbol{x}+e_{j})-g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$, where $e_{j}$ is the unit vector indicating $D_{j}$ is selected and $p_{j}$ is set as zero, and
  2. 2.
    Budget surplus based on the prior decision: $\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1}):=b_{k}-(\tilde{\boldsymbol{x}}^{k,N-1})^{T}\boldsymbol{p}$.
A buyer’s maximum willingness to pay (MWP) is defined as the highest price buyer $B_{k}$ is willing to pay – based on marginal utility – and able to pay – based on budget surplus:
$\text{MWP}_{k}:=\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}\{\min\{\Delta u_{k}(\boldsymbol{x})^{+},\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1})\}\},$ (9)
where $\Delta u_{k}(\boldsymbol{x})^{+}:=\{\Delta u_{k}(\boldsymbol{x}),0\}$ denotes the positive part of marginal utility, ensuring buyer $B_{k}$’s $\text{MWP}_{k}$ is non-negative – if the marginal utility of $D_{j}$ is negative, the buyer will not purchase it.
Then the seller’s optimal price $p_{j}^{*}$ is the MWP that results the largest profit across buyers:
Lemma 2 (Characterization of Optimal Price $p_{j}^{*}$).
Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP (see Appendix˜D for proof):
$p_{j}^{*}\in\cup_{k=1}^{M}\text{MWP}_{k}.$ (10)
We see that optimal price $p_{j}^{*}$ is fairshare for the seller: it maximizes seller’s profit while aligning with dataset’s utility and buyer’s budget.
As shown in Figures˜2(c) and 2(d), we simulate a data market with $2$ sellers and $2$ buyers, where players make one-shot decisions. Seller $S_{1}$ sets its price first, followed by seller $S_{2}$. Each plot shows each seller’s profit function $r(p_{j})$, with breakpoints at each buyer’s MWP. Pricing above a buyer’s MWP leads to a drop in sales. Thus, the seller’s optimal price aligns with the MWP that yields maximum revenue.
Buyer-Side: Fairshare Is Overall Optimal. We now show that the fairshare price $p_{j}^{*}$, derived from seller optimization, is also optimal from the buyer’s perspective. As established in Section˜3.2, exploitative pricing leads to persistent seller exit. In contrast, fairshare pricing ensures full participation and maximizing the buyer’s overall utility. To formalize this, we consider the same infinite-horizon setting in Section˜3.2 with a single buyer and seller. Notably, the fairshare price is the ideal price for sellers as it maximizes its profit. Therefore, under fairshare pricing, the sellers will maintain sustained participation. Then, we obtain:
Lemma 3 (The Optimal Price for Buyer Is Also $p^{*}_{t}$).
The fairshare price for the seller $S$ under LLM data framework is $p^{*}_{t}=\min\{u_{t},b_{t}\},\forall t$. With Assumptions˜1, 1.1 and 2, $p^{*}_{t}$ is the optimal pricing strategy that maximizes buyer’s expected cumulative utility (Equation˜8).
Figure 3: Analysis of the buyer’s cumulative net utility as a function of the acquisition prices over time ($T=1,2,5,10$). Note: the market has one buyer and seller.
In Figure˜3, we develop a synthetic simulation illustrating that a buyer’s optimal pricing strategy rapidly converges to the fairshare price. Consider a simulation with players receive constant utility $u_{t}=2$ and budget $b_{t}=1,\forall t$, with the discount factor $\delta=0.95$, so the seller’s optimal price remains $p^{*}_{t}=1$. At each time step, the buyer selects a price that maximizes cumulative net utility. The buyer begins with a low, exploitative price at $T=1$, but by $T=5$, converges to the fairshare price and maintains it thereafter. This illustrates our core theoretical insight: fair pricing emerges as the optimal long-run strategy when buyers account for overall market sustainability.
Next, we also explore the role of the discount factor $\delta$ in our framework:
Lemma 4 (The Trade-Off Threshold Is Increasing as $\delta$ Decreases).
The threshold time period where the fairshare pricing obtains higher cumulative utility than any class of exploitative pricing is:
$t^{*}:=\sup_{p_{t}<p_{t}^{*},\forall t}\left\{T\in\mathbb{N}:\mathbb{E}\left[\sum_{t=0}^{T}\delta^{t}\left(\left(u_{t}-p^{*}_{t}\right)-\prod_{i=0}^{t-1}\pi(p_{i},p^{*}_{i})\left(u_{i}-p_{i}\right)\right)\right]\leq 0\right\}.$ (11)
And $t^{*}$ is increasing as $\delta$ decreases.
We run detailed robustness check with different values of $\delta$ for our experiments in Section˜E.2.
4 Empirical Analyses: Benefits of Fairshare Pricing
This section evaluates the LLM data market and the proposed fairshare pricing framework.
4.1 Experimental Setup
In our experiments, a buyer is equipped with a single LLM, and each seller owns one data sample. Buyers seek to buy training data to improves task-specific model performance (e.g.,math problem solving, medical diagnosis, or physical reasoning), which in turn increases their utility. (We assume an affine mapping between performance and utility; see Section˜C.1.)
Buyers and Models. We consider three buyers, each using a standard open-source LLMs: Llama-3.2-Instruct-1b [grattafiori2024llama3herdmodels], Pythia-1b, and Pythia-410m [biderman2023pythia]. These models are pre-trained on different corpora and exhibit varying preferences for downstream post-training data [mai2024fine].
Sellers and Datasets. We focus on challenging, human-annotated tasks: MathQA and GSM8K [amini2019mathqa, cobbe2021training] for math, MedQA [jin2020medqa] for medical diagnosis, and PIQA [bisk2020piqa] for physical reasoning [lu2022survey, liu2023mathematical, ahn2024large]. Table˜1 in Appendix˜F shows dataset splits and examples. We use the training splits as seller data and simulate the market dynamics from Section˜3, treating each task as a market scenario.
4.2 LLM Data Market Experiments
Figure 4: (1) buyer’s cumulative utilities with high- (Figure˜4(a)) and low-budget buyer (Figure˜4(b)), and (2) sellers’ average cumulative profits (Figure˜4(c)) and active seller numbers (Figure˜4(d)) over $100$ time periods. Pythia-1b; MedQA; Groups: (1) fairshare, (2) reduced, (3) random, and (4) exploitative.
We first evaluate our pricing framework in terms of buyer and seller welfare.
Market Design: Following the general setups in Section˜4.1, we simulate a market with $2$ buyers and $10$ sellers over multiple time steps. To examine buyers with varying resources, we include a high-budget buyer (well-funded LLM builder) and a low-budget buyer (under-resourced one). Each buyer’s budget is randomly drawn from distributions with different mean. At each time step, (1) sellers arrive sequentially with a new dataset ( $300$ data samples) at a fixed price, and then (2) once all sellers arrive, buyers make purchases by solving Equation˜2. Full details are in Section˜E.2.
Participation Function: Following Assumptions˜1, 1.1 and 2, we simulate $100$ time steps with a discount factor $\delta=0.98$. We also run experiments with different value of $\delta$ showing consistent and robust results. (See Appendix˜E). Seller participation probability is defined as $\pi(p_{j,t},p^{*}_{j,t})=p_{j,t}/p^{*}_{j,t}$ for its simplicity and compliance with Assumptions˜1 and 1.1. Sellers receiving exploitative pricing ( $p_{j,t}<p^{*}_{j,t}$) are less likely to participate.
[…]
Figure˜4 compares the overall welfare outcomes of different pricing methods for buyers and sellers using Pythia-1b on the MedQA task. Results on MathQA and PIQA (with Pythia-410m and Llama-3.2-Instruct-1b) in Appendix˜F show similar patterns.
(a) High-budget buyer.
(b) Low-budget buyer.
Figure 5: Purchased datasets for the buyer with high budget (Figure˜5(a)) and low budget (Figure˜5(b)) over $100$ time periods. Pythia-1b; MedQA.
Exploitative Pricing Leads to Lose-Lose Outcomes: The exploitative pricing method, which sets uniformly low, fixed prices to reflect real-world practices ([cbsnews2024]), offers LLM developers short-term utility gains (see Figures˜4(a) and 4(b)). However, this approach systematically undervalues datasets and unfairly compensates annotators. As a result, sellers exit the market over time, triggering a collapse in data supply. Even well-funded LLM developers are unable to source sufficient data – hindering the advancement of LLMs.
Fairshare Pricing Leads to Win-Win Outcomes: For sellers, fairshare pricing consistently yields the highest profit over time (Figure˜4(c)), aligning with predictions from our theoretical model (Section˜3.1). For buyers, fairshare pricing framework is particularly effective for the high-budget buyer (Figure˜4(a)), maximizing long-term utility. However, the low-budget buyer (Figure˜4(b)) experiences reduced short-term utility in exchange for long-term gains. Its limited budget prevents them from fully leveraging the increased dataset supply ensured by fairshare pricing, making other low-pricing methods initially more appealing. Yet, in the long run, fairshare pricing sustains seller participation, ensuring data supply.
4.3 Ablation Study: Effect of Different Data Valuation Methods
This experiment assesses the impact of different data valuation strategies in the fairshare framework.
Setup. Using the models and datasets introduced earlier in Section˜4.1, we run separate simulations for each market (math, medical, physical reasoning), testing different data valuation methods for the buyer. Each buyer is randomly assigned to use one of four valuation methods to select training data and fine-tune their model. The seller receives payments according to the data’s assigned value. Valuation scores are normalized to $[0,1]$ for pricing compatibility.
Data Valuation Methods. We consider the following methods, where each assigns a value to every data sample: (1) Constant baseline – assigns the same value, mimicking flat-rate pricing on platforms; (2) Random baseline – values drawn uniformly from $[0,1]$; (3) Semantic – uses BM25 [trotman2014improvements] to compute average similarity to the representative set; (4) Influence-based: returns a score which leverages learning gradients to estimate a data sample’s avg. contribution to model learning of a representative dataset. Specifically, we use $\text{Infl}_{\text{IP}}$ [pruthi2020tracin, xia2024less] and DataInf  [kwon2023datainf], which are influence-based methods adapted for the LLM realm. See Appendix˜A for additional details.
Furthermore, we adopt a one-step training approach for data valuation, where the value of each data sample is estimated by performing a single training step on it and measuring the resulting change in model performance relative to the original model. Results are presented in in Section˜F.2. This approach serves as an “oracle” for influence-based methods [yu2024mates, pruthi2020tracin], as it directly quantifies data value through model training, although it is more computationally expensive than the previous listed methods.
Market/Pricing Setup: We reserve $1\%$ of the samples from each dataset’s training split to represent the existing data in their respective markets. Each data sample was randomly priced between $(0,1]$. Next, for each remaining data sample in the training set, we determine whether each buyer will purchase the data sample at potential price points $[0.5,0.625,0.75,0.875,1.0]$ by solving Equation˜2. The seller then sets their prices according to Equation˜4. We price data separately for each data valuation method. This assesses the method’s ability to discern whether a new data sample is worth purchasing for each buyer given the existing market data, as noted in our analysis in Section˜3. Further details on all experiments are described in Section˜E.1.
Results. Figure˜6 presents results across data valuation methods. Buyers using a valuation method (ie., BM25, $\text{Infl}_{\text{IP}}$, DataInf) in general achieved higher model performance across tasks. When considering the trade-off between cost and performance, $\text{Infl}_{\text{IP}}$ offered the best balance, delivering strong model improvements at a lower cost than constant, random, and BM25 (Figures 6(a) and 9(a)). Our results highlight the benefits of learning-aware data valuation methods. By prioritizing high-impact data, they offer a better alternative for buyers, particularly those with limited budgets.
We also run the error analysis of data valuation methods. Following previous studies [jiao2024context, wang2024greats], we compute the correlation between Oracle and $\text{Infl}_{\text{IP}}$, reporting spearman correlation of $0.54,0.42$ for MathQA and PIQA (see Figure˜8 in Appendix˜F). This underscores the opportunity for future advances in data valuation accuracy and scalability, which can be seamlessly integrated into our flexible fairshare framework.
Figure 6: Left column: Average price-per-sample cost of purchased data across math, medical, and physical reasoning data markets using different data valuation methods. Right, middle-columns: Buyers’ model performance versus cost, before purchasing (BP) data, and after purchasing 2K and 4K data samples. Additional analysis on the Pythia-1b/410m models are in Appendix˜F.
5 Conclusion
We introduced a fairshare pricing framework that leverages data valuation methods to enable transparent and equitable training data pricing for LLMs. Our results show that buyers achieve higher gains at lower cost. At the same time, sellers earn optimal prices for their data, fostering a win-win outcome that enhances long-term market sustainability and social welfare. To our knowledge, this is the first work to integrate game-theoretic pricing models with LLM-specific data valuation techniques, addressing the real-world dynamics of the emerging data market. Our framework offers actionable insights for policymakers and regulators aiming to ensure fairness and transparency in LLM training data markets. By fostering fair market access, our framework also empowers small businesses and startups, leading to more equitable technological advancements. Future work could extend our framework to incorporate additional data valuation methods, incomplete information settings (e.g., Bayesian games), and diverse data domains (e.g., pretraining vs. fine-tuning). We hope this work paves a fruitful way for future research in equitable markets for AI and emerging technologies.
Appendix A Influence-based Data Valuation
In Section 4, we introduced a gradient-based data attribution method, denoted as $\text{Infl}_{\text{IP}}$. In this section, we provide additional information on $\text{Infl}_{\text{IP}}$, which has been shown to be effective in training data selection in previous works [pruthi2020tracin, xia2024less]. Suppose we have a LLM parameterized by $\theta$, and a train set $D$ and a test set $\mathcal{D}^{\prime}$. For a training sample $d\in D$, we wish to estimate its training impact on a test sample $d^{\prime}\in D$. That is, we want to measure the impact of $d$ on the model’s loss on $d^{\prime}$ (i.e., $\mathcal{L}(d^{\prime};\theta)$). As simple method of achieving this is to take training step – that is, a gradient descent step – on $d$ and obtain:
$\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$ (12)
where $\eta$ is the learning rate. Then, in order to measure the influence of $d$ towards $d^{\prime}$, we wish to find the change in loss on $d^{\prime}$:
[…]
Instead of taking a single training step to measure the influence of $d\in D$ on $d^{\prime}$, we can instead approximate Equation˜13 with using the following:
Lemma 5.
Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
$\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})\approx\nabla\mathcal{L}(d^{\prime};\theta)\cdot\nabla\mathcal{L}(d;\theta)$
See Appendix˜D for the proof.
Then, we set $\text{Infl}_{\text{IP}}$ to be:
$\text{Infl}_{\text{IP}}=\nabla\mathcal{L}(d^{\prime};\theta)\cdot\nabla\mathcal{L}(d;\theta)$ (14)
which is the dot-product between the learning gradients of $d^{\prime}$ and $d$.
Appendix B Royalty model
So far, we have shown the case of flat rate (see Section˜3.1), which is well-suited resource-rich buyers, such as leading tech companies whose LLMs generate significant economic value due to their wide-ranging impact and scalability. In this section, we introduce the royalty model, a contract framework that differs from the flat rate by offering a subscription-like structure. Under the royalty model, the price paid for training data is proportional to the future economic value generated by the LLM, providing a flexible and performance-based approach to data valuation. This scenario incorporates buyers in a less dominant position – those who are (1) uncertain about the prospective model outcome or (2) do not own a sufficient cash flow for purchasing data with full prices. We present updated decision-making models for buyers and sellers as follows.
Buyers. Unlike the flat pricing setting, the buyer $B_{k}$ would alternatively pay with a fractional price. Suppose each dataset $D_{j}$ is priced with an individual rate $\alpha_{j}\in[0,1)$ (as we denote $\boldsymbol{\alpha}=(\alpha_{1},\cdots,\alpha_{N})$), then the price of an arbitrary data collection $u_{k}\left(\boldsymbol{x}\right)$ is a fraction of its future marginal gain, i.e., $\boldsymbol{x}^{T}\boldsymbol{p}=f(\boldsymbol{\alpha},\boldsymbol{x})u_{k}\left(\boldsymbol{x}\right)$, where the overall rate function $f:[0,1)^{|\boldsymbol{\alpha}|}\times\{0,1\}^{|\boldsymbol{x}|}\rightarrow[0,1)$ depends on specific contexts. We assume that $f$ is a monotonically non-decreasing function of $\boldsymbol{\alpha}$. In this sense, the buyer $B_{k}$ reduces the risk of losing $b_{k}$ from its cash flow while the seller is betting on the potential value of the LLM $M_{k}$. Then we obtain an updated objective function for $B_{k}$:
$g_{k,N,\text{frac}}(\boldsymbol{x}):=(1-f(\boldsymbol{\alpha},\boldsymbol{x}))u_{k}\left(\boldsymbol{x}\right),$ (15)
On the other hand, similar to the budget constraint (see Equation˜2), here each buyer $B_{k}$ has a maximum rate $\overline{\alpha}_{k}$ that it is willing to pay. Then the buyer’s purchasing problem is given as
$\tilde{\boldsymbol{x}}^{k,N,\text{frac}}$ $:=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}_{k,N,\text{frac}}}\;g_{k,N,\text{frac}}(\boldsymbol{x}),\quad\text{s.t.}$ (16)
$\mathcal{X}_{k,N,\text{frac}}$ $:=\{\boldsymbol{x}\mid g_{k,N,\text{frac}}(\boldsymbol{x})\geq 0,f(\boldsymbol{\alpha},\boldsymbol{x})\leq\overline{\alpha}_{k}\},$ (17)
And $\tilde{\boldsymbol{x}}^{k,N,\text{frac}}$ is the optimal solution to $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N,\text{frac}}}g_{k,N,\text{frac}}(\boldsymbol{x})$ with a given rate vector $\boldsymbol{\alpha}$.
Sellers. In the fractional pricing setting, since the buyer $B_{k}$ pays for the entire data collection, there should exists a fair and transparent allocation mechanism that distributes a portion of the total price charged to each individual dataset $D_{j}$. That is, $\boldsymbol{x}_{j}p_{j}=\sum_{k=1}^{M}f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N,\text{frac}})u_{k}(\tilde{\boldsymbol{x}}^{k,N,\text{frac}})$, where $f(\cdot)=\sum_{j=1}^{N}f_{j}(\cdot)$. And we assume that for all $j\in[N]$, $f_{j}(\cdot)$ is monotonically non-decreasing over $\boldsymbol{\alpha}$. Therefore, we have an updated profit function for $Se_{j}$:
$r_{\text{frac}}(\alpha_{j})$ $:=\sum_{k=1}^{M}f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N,\text{frac}})u_{k}(\tilde{\boldsymbol{x}}^{k,N,\text{frac}})-c_{j}.$ (18)
which gives the following problem:
$\alpha^{*}_{j}$ $:=\operatorname*{arg\,max}_{{\alpha}_{j}\in\mathcal{A}_{j,M,\text{frac}}}\;r_{\text{frac}}(\alpha_{j}),\quad\text{s.t.}$ (19)
$\mathcal{A}_{j,M}$ $:=\{{\alpha}_{j}\in[0,1)\mid r_{\text{frac}}(\alpha_{j})\geq 0\},$ (20)
From this point onward, the market dynamics stays the same as in the previous section. It is noted that, compared to $\max_{{p}_{j}\in\mathcal{P}_{j,M}}r(p_{j})$ where optimal flat rate is indirectly connected to the utility, the optimal rate of $\max_{{\alpha}_{j}\in\mathcal{A}_{j,M,\text{frac}}}r_{\text{frac}}(\alpha_{j})$ offers a more direct representation of the utility.
B.1 Solving for optimal price of the royalty model
Similar to flat rate, in the case of royalty model, we need to solve buyer’s problem $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}}g_{k,N-1,\text{frac}}(\boldsymbol{x})$ (before the arrival of $S_{j}$) and $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N,\text{frac}}}g_{k,N,\text{frac}}(\boldsymbol{x})$ (after the arrival of $S_{j}$) for all $k\in[M]$ and seller’s problem $\max_{{\alpha}_{j}\in\mathcal{A}_{j,M,\text{frac}}}r_{\text{frac}}(\alpha_{j})$.
Solve buyer’s problems. For each feasible collection of datasets $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}$ (before the arrival of $S_{j}$), denotes it union with dataset $D_{j}$ as $\boldsymbol{x}^{\text{new}}$. Then we run the check: (1) if the net utility of $\boldsymbol{x}^{\text{new}}$ is larger than the one of $\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}$, i.e., $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$, and (2) if the rate for purchasing $\boldsymbol{x}^{\text{new}}$ is still under the budget $\overline{\alpha}_{k}$, i.e., $f(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}$, where $\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix}$ denotes concatenating $\alpha_{j}$ to $\boldsymbol{\alpha}$. If the answer is positive to bother tests, then we can determine that the buyer $B_{k}$ will change its decision and purchase $S_{j}$ under the rate $\alpha_{j}$.
Solve seller’s problem. First, we consider when $\alpha_{j}=0$. We could first find all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$. And we denote the set that contains such $\boldsymbol{x}$ as $\mathcal{X}_{k,N-1,\text{frac}}^{1}$. If $\mathcal{X}_{k,N-1,\text{frac}}^{1}$ is empty, then $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}=0$ (indicating whether $B_{k}$ will purchase $D_{j}$ at price $p_{j}$ or not), as $S_{j}$ cannot bring positive value to $B_{k}$; else, then for all $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, thanks to the monotonicity of $f_{j}$ over $\alpha_{j}$, we could gradually increase $\alpha_{j}$ until the either of the two criterion are met first: (1) we find the largest $\alpha_{j}$ such that $g_{k,N,\text{frac}}(\boldsymbol{x}^{\text{new}})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$, and (2) $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}$. Then we have the following property about the optimal rate $\alpha_{j}^{*}$ for Equation˜19:
Lemma 6 (Characterization of $\alpha_{j}^{*}$ under royalty model).
Define $\alpha_{j}^{\boldsymbol{x}}$ as
$\min$ $\left\{\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})}{u_{k}(\boldsymbol{x}^{\text{new}})}\right\},\right$.
$\left.\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}\right\}\right\}.$ (21)
For every $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$ and all $k\in[M]$, we obtain $\alpha_{j}^{\boldsymbol{x}}$ and their union $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. Then we have $\alpha_{j}^{*}\in\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$.
Remark B.1 (Similarities between flat rate and royalty model).
Observing from Lemmas˜6 and 2, we see that the both the optimal price $p^{*}_{j}$ and the optimal rate $\alpha_{j}^{*}$ are closely tied to $B_{k}$’s maximum willingness to pay. That is, compared to the market prior to the arrival of $S_{j}$, the optimal values are characterized by the minimum of two factors: (1) marginal utility that $S_{j}$ provides to $B_{k}$ and (2) $B_{k}$’s budget surplus. It is also noted that, under royalty model, the rate function $f$ also plays an important role as it determines the how the single rate $\alpha_{j}$ affects the total rate that $B_{k}$ pays.
Appendix C Applications for Real-Life Scenarios
In real-life settings, the relationship between the data valuation of a training sample and the buyer’s utility $u_{k}$ (i.e., the economical value, which may be expressed in dollar amounts) can have different mappings, as mentioned in Section˜3.1. Suppose the data valuation function is denoted as $v_{k}:D\rightarrow\mathbb{R}$ for a dataset $D$. Then, a buyer may expect a linear relationship between $v_{k}$ and $u_{k}$, where the utility increases as the data valuation score increases. Alternatively, a buyer may prefer to only purchase data beyond a certain threshold for $v_{k}$. In this section, we present three types mappings between $v_{k}$ and $u_{k}$ to reflect these scenarios: linear, discrete, and zero-one mappings. We show that these mappings can be easily adapted to our proposed framework in Section˜3. We only present the updated buyer’s purchasing problem (Equation˜2) since the seller’s pricing problem (Equation˜4) stays the same.
C.1 Linear Outcome
In practice, there are many applications where $u_{k}$ is an affine function of $v_{k}$. As previously mentioned, training LLMs on data with higher valuation scores $v_{k}$ can result in better economic value towards downstream model performance, as shown in previous works [xia2024less, yu2024mates]. In this outcome setting, in addition to considering $u_{k}$ to be a affine function of $v_{k}$, we also include a bias variable $\beta$ to account for other potential other factors that are independent of $v_{k}$. Therefore, we can set $u_{k}=\gamma v_{k}(\boldsymbol{x})+\beta$ into Equation˜1, where $\gamma\in\mathbb{R}_{+}$ is a known coefficient, and obtain buyer $B_{k}$’s net utility function for the linear outcome:
$g_{k,N}(\boldsymbol{x})=\gamma v_{k}(\boldsymbol{x})+\beta-\boldsymbol{x}^{T}\boldsymbol{p},$ (22)
To obtain optimal price $p^{*}$, we can directly refer to same procedure described in Section˜3 using set values for $\gamma$ and $\beta$.
C.2 Discrete Outcome
There are also many applications where $u_{k}$ is discrete. For instance, if the data buyers are participating in an LLM benchmark challenge, such as MMLU [hendrycks2021mmlu], then training on data that falls within various ranges $v_{k}$ may lead to drastically different model performance, and hence leaderboard rankings.
To mirror this, consider $u_{k}$ to be a category variable. We denote $\{c_{h}\}_{h=1}^{H}$ as a strictly increasing set of numbers such that when $v_{k}\in[c_{h},c_{h+1})$, the buyer will receive reward $u_{k,h}$. We also assume that $u_{k,h+1}>u_{k,h}$ since higher data valuation scores may lead to a larger reward. Therefore, we could set $u_{k}=\sum_{h=1}^{H}\mathbbm{1}_{\{v_{k}(\boldsymbol{x})\in[c_{h},c_{h+1})\}}u_{k,h}(\boldsymbol{x})$ and rewrite buyer $B_{k}$’s net utility function as
$g_{k,N}(\boldsymbol{x})=\sum_{h=1}^{H}\mathbbm{1}_{\{v_{k}(\boldsymbol{x})\in[c_{h},c_{h+1})\}}u_{k,h}(\boldsymbol{x})-\boldsymbol{x}^{T}\boldsymbol{p}.$ (23)
We again apply the same procedure in Section˜3 to solve for the optimal pricing.
C.3 Zero-One Outcome
There are scenarios where the data buyers are risk-adverse and focus on the effects of rare events. In these cases, suppose that $v_{k}$ is normalized between [0, 1]. Then buyers may wish to purchase training data with higher values of $v_{k}$, assuming that purchasing data with lower $v_{k}$ may result in severe adverse effects. For instance, data buyers who are building AI for healthcare should not purchase data with incorrect medical information, and even a small amount of contaminated data can result in severe real-life consequences such as mis-diagnosis [jin2020medqa, zhou2023survey] or unsuitable medical protocols in emergency situations [sun2024edcopilotreduceemergencydepartment]. Therefore, in this context, we consider $u_{k}$ as a generalized Bernoulli distribution. The downstream outcome has a small positive reward $\underline{u}$ with probability $v_{k}$ (normal events) and a massive negative reward $\overline{u}$ with probability $1-v_{k}$ (undesirable rare events). And we assume that $\mathbb{E}(u_{k})>0$. Therefore, we can plug in and obtain buyer $B_{k}$’s net utility function:
$g_{k,N}(\boldsymbol{x})$ $=\mathbb{E}[u_{k}(\boldsymbol{x})]-\boldsymbol{x}^{T}\boldsymbol{p}$ (24)
$=v_{k}(\boldsymbol{x})(\underline{u}-\overline{u})+\overline{u}-\boldsymbol{x}^{T}\boldsymbol{p},$ (25)
which is an affine function of $v_{k}$. Therefore, we again apply same procedure in Section˜3 to solve for the optimal pricing.
C.4 Multiple tasks
In practice, many LLMs are evaluated over multiple tasks [hendrycks2021mmlu]. To this end, we consider the context where buyer $B_{k}$ wishes their model $\mathcal{M}_{k}$ to perform well across multiple tasks, denoted as $Q$. Each data valuation score for a task is denoted by $v^{k}_{1},\cdots,v^{k}_{Q}$ and the vector of all task valuations is denoted as $\boldsymbol{v_{k}}=(v_{k,1}\cdots v_{k,Q})$. Then we consider that the utility $u_{k}$ is an affine function of the utility in each task, denoted by $\boldsymbol{u_{k}}=(u_{k,1}\cdots u_{k,Q})$ that is, $u_{k}=\boldsymbol{\theta}^{T}\boldsymbol{u_{k}}+\epsilon$, where $\boldsymbol{\theta}\in\mathbb{R}^{Q}$ is a coefficient vector and $\epsilon\in\mathbb{R}$ denotes other factors independent from $\boldsymbol{u_{k}}$. We also assume that the each task is one of three categories mentioned in the last section. Therefore, we can rewrite $\boldsymbol{u_{k}}$ as a function of $\boldsymbol{v_{k}}$, which gives $u_{k}=\boldsymbol{\theta}^{T}\boldsymbol{u_{k}}(\boldsymbol{v_{k}})+\epsilon$. Therefore, the buyer’s net utility function becomes
$g_{k,N}(\boldsymbol{x})=\boldsymbol{\theta}^{T}\boldsymbol{u_{k}}(\boldsymbol{v_{k}}(\boldsymbol{x}))+\epsilon-\boldsymbol{x}^{T}\boldsymbol{p}.$ (26)
whose solution could adopt the same procedure as described in Section˜3 to solve for the optimal pricing.
Appendix D Lemmas and Proofs
Lemma˜1. With Assumptions˜1, 1.1 and 2, any exploitative pricing (i.e., $p_{t}<p_{t}^{*},\forall t$) will only maximize cumulative utility within a finite horizon – after which it is strictly suboptimal.
Proof.
Lemma˜1 equivalently states that with Assumptions˜1, 1.1 and 2, the optimal pricing strategy for the buyer is also the ideal price $p^{*}_{t}$. Thus, we show that when the buyer pays the ideal price $p^{*}_{t}$, its total value is the largest, i.e.,
$\mathbb{E}\left[u_{t}-p^{*}_{t}+\delta\mathbb{E}\left[r(p^{*}_{t},p^{*}_{t})G\mid u_{t},b_{t}\right]\right]\geq\mathbb{E}\left[u_{t}-p_{t}+\delta\mathbb{E}\left[\pi(p_{t},p^{*}_{t})G\mid u_{t},b_{t}\right]\right]$ (27)
for all $p_{t}\in[0,p^{*}_{t})$.
The seller will not set a price above its ideal price $p^{*}_{t}$ as it will decreases its profit, since the ideal price $p^{*}_{t}$ should be its profit-maximizing price. Any $p_{t}>p^{*}_{t}$ results lower profits.
Also, the buyer will not accept a price $p_{t}>p^{*}_{t}$, since it decreases the net utility gain $u_{t}-p_{t}$ without increasing seller’s participation probability $\prod_{t=0}^{T-1}\pi(p_{t},p^{*}_{t})$.
Through some linear transformation, this is equivalent to show that
$\mathbb{E}\left[p_{t}-p^{*}_{t}+\delta\mathbb{E}\left[G(\pi(p^{*}_{t},p^{*}_{t})-\pi(p_{t},p^{*}_{t}))\mid u_{t},b_{t}\right]\right]\geq 0.$ (28)
We first find the lower bound of $G$. We see that $p^{*}_{t}$ gives a payoff of
$\mathbb{E}\left[\sum_{t=0}^{\infty}\delta^{t}\left(u_{t}-p^{*}_{t}\right)\right]\geq\frac{\min_{t\in[0,\infty)}\mathbb{E}[u_{t}-p^{*}_{t}]}{1-\delta}.$ (29)
Therefore, we must have $G\geq\frac{\min_{t\in[0,\infty)}\mathbb{E}[u_{t}-p^{*}_{t}]}{1-\delta}$. Along with Assumptions˜1, 1.1 and 2, this gives us, for a given $u_{t}$ and $b_{t}$,
$\frac{\delta G\left(\pi(p^{*}_{t},p^{*}_{t})-\pi(p_{t},p^{*}_{t})\right)}{p^{*}_{t}-p_{t}}\geq\delta GL\geq 1,$ (30)
implying that
$\mathbb{E}\left[p_{t}-p^{*}_{t}+\delta\mathbb{E}\left[G(\pi(p^{*}_{t},p^{*}_{t})-\pi(p_{t},p^{*}_{t}))\mid u_{t},b_{t}\right]\right]\geq 0.$ (31)
Lemma 7 (Participation Loss Is Lower-Bounded).
With Assumption˜1, let $P:=\lim_{T\rightarrow\infty}P_{T}$ and $S:=\sum_{i=0}^{\infty}(1-\pi(p_{i},p^{*}_{i}))$. Then for the class of all exploitative pricing strategies where $p_{t}\leq p_{t}^{*},\forall t$, we have (1) $P_{T}$ is strictly decreasing and (2) The limit of reduced participation is lower bounded, i.e., $1-P\geq 1-e^{-S}>0$.
Proof.
The first part is trivial to see as we assume that $\pi(p_{i},p^{*}_{i})<1$ for all $p_{i}<p^{*}_{i}$ from Assumption˜1.
For the second part, we first denote $P:=\lim_{t\rightarrow\infty}P_{t}$. Then we take the following transformation:
$\log P=\log\left(\prod_{i=0}^{\infty}\pi(p_{i},p^{*}_{i})\right)=\sum_{i=0}^{\infty}\log\left(\pi(p_{i},p^{*}_{i})\right).$ (32)
Since we have $0<\pi(p_{i},p^{*}_{i})<1$, then $\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\pi(p_{i},p^{*}_{i})-1$, indicating that
$\log P=\sum_{i=0}^{\infty}\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\sum_{i=0}^{\infty}\left(\pi(p_{i},p^{*}_{i})-1\right)=-S,$ (33)
which gives $P\leq e^{-S}$ by exponentiating both sides. ∎
Lemma˜2. Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP:
$p_{j}^{*}\in\cup_{k=1}^{M}\text{MWP}_{k}.$ (34)
Proof.
Recall that without dataset $D_{j}$, each buyer $B_{k}$ has already solved $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}g_{k,N-1}(\boldsymbol{x})$ according to our market definition in Section˜3, where $\mathcal{X}_{k,N-1}$ is the set of all feasible purchase decisions. Next, after seller $S_{j}$ (with $D_{j}$) has arrived on the market, we analyze the conditions in which $B_{k}$ will purchase $D_{j}$ at a potential price $p_{j}$. For each feasible purchase decision (i.e., a collection of datasets), represented by $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$, let $\boldsymbol{x}+e_{j}$ denote its union with $D_{j}$, where $e_{j}$ is the unit vector indicating $D_{j}$ is selected. For buyer $B_{k}$ to change their previous decision to purchase $D_{j}$, there are two requirements that need to be satisfied. First, we must have:
$g_{k,N}(\boldsymbol{x}+e_{j})>g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1}).$ (35)
That is, the net utility $g_{k,N}(\boldsymbol{x}+e_{j})$ of purchasing decisions $\boldsymbol{x}+e_{j}$, must be larger than the net utility $g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$ of a previous optimal purchasing decision $\tilde{\boldsymbol{x}}^{k,N-1}$. It is also noted that $g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})=g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$. Second, for buyer $B_{k}$ to purchase $\boldsymbol{x}+e_{j}$ at price $p_{j}$, we must fulfill the budget constraint:
$p_{j}\leq b_{k}-\left(\tilde{\boldsymbol{x}}^{k,N-1}\right)^{T}\boldsymbol{p}=\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1}),$ (36)
which ensures that purchasing $D_{j}$ does not exceed the buyer’s budget $b_{k}$. If both requirements are satisfied, then the buyer $B_{k}$ will change their previous purchasing decision in order to purchase $D_{j}$ under the price $p_{j}$. This procedure is presented in detail in Algorithm˜1 in Section˜F.1.
Next, given the conditions for the buyer $B_{k}$ to purchase $D_{j}$, the seller must solve $\max_{{p}_{j}\in\mathcal{P}_{j,M}}r(p_{j})$ to find the optimal price $p_{j}^{*}$. First, we consider an edge case where the price of dataset $D_{j}$ is set as $p_{j}=0$. For a buyer $B_{k}$, we denote $\mathcal{X}_{k,N}^{1}$ as the set of all purchasing decisions where including $D_{j}$ in the purchase improves the buyer’s previous net utility $g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})$. That is, for every $\boldsymbol{x}+e_{j}\in\mathcal{X}_{k,N}^{1}$, we have $g_{k,N}(\boldsymbol{x}+e_{j})>g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})$. If $\mathcal{X}_{k,N}^{1}$ is empty, then $B_{k}$ will not purchase $D_{j}$ at any price, since $D_{j}$ cannot bring positive improved net utility to $B_{k}$. Then, when $p_{j}$ gradually increases and exceeds $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}\{\min\{\Delta u_{k}(\boldsymbol{x}+e_{j}),\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1})\}\}$, then $B_{k}$ will decide not to purchase $D_{j}$, causing the value of $\sum_{k=1}^{M}\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})$ to drop by one. Since the profit function $r(p_{j})$ is a piecewise linear function, the its optimal point must be one of its breakpoints. ∎
Lemma˜3 The optimal price for the seller $S$ under our framework is
$p^{*}_{t}=\min\{u_{t},b_{t}\},\forall t.$ (37)
With assumptions˜1, 1.1 and 2, $p^{*}_{t}$ gives the buyer the maximum cumulative net utility over infinite horizon.
Proof.
In a single buyer and seller setting, we could trivially see that the optimal price for the seller $p_{t}^{*}=\min\{u_{t},b_{t}\}$: if $p_{t}>p_{t}^{*}$, then the buyer would not purchase this dataset since the net utility would be negative; if $p_{t}<p_{t}^{*}$, then the seller’s profit is not maximized.
Further, we refer to the proof of Lemma˜1 for the second part. ∎
Lemma˜4 The threshold time period where the fairshare pricing obtains higher cumulative utility than any class of exploitative pricing is:
$t^{*}:=\sup_{p_{t}<p_{t}^{*},\forall t}\left\{T\in\mathbb{N}:\mathbb{E}\left[\sum_{t=0}^{T}\delta^{t}\left(\left(u_{t}-p^{*}_{t}\right)-\prod_{i=0}^{t-1}\pi(p_{i},p^{*}_{i})\left(u_{i}-p_{i}\right)\right)\right]\leq 0\right\}.$ (38)
And $t^{*}$ is increasing as $\delta$ increases.
Proof.
For any given class of exploitative pricing strategy, when $\delta$ increases, the part: $\sum_{t=0}^{T}\delta^{t}\left(\left(u_{t}-p^{*}_{t}\right)-\prod_{i=0}^{t-1}\pi(p_{i},p^{*}_{i})\left(u_{i}-p_{i}\right)\right)$ increases. Therefore, for all class of exploitative pricing strategy, i.e., $p_{t}<p_{t}^{*},\forall t$, then $t^{*}$ also decreases. ∎
Lemma˜5. Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
$\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})\approx\nabla\mathcal{L}(d^{\prime};\theta)\cdot\nabla\mathcal{L}(d;\theta)$
Proof.
First, we consider the change in loss of $z^{\prime}$ using a first-order approximation:
[…]
Next, suppose a gradient descent step is taken on training sample $d$, and the model parameters are updated as: $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Thus, we have $\hat{\theta}-\theta=-\eta\nabla\mathcal{L}(d;\theta)$, and the change in loss can be written as
[…]
Given that $\eta$ is a constant. ∎
Lemma˜6 Define $\alpha_{j}^{\boldsymbol{x}}$ as
$\min$ $\left\{\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})}{u_{k}(\boldsymbol{x}^{\text{new}})}\right\},\right$.
$\left.\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})\leq\overline{\alpha}_{k}\right\}\right\}.$ (42)
For every $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$ and all $k\in[M]$, we obtain $\alpha_{j}^{\boldsymbol{x}}$ and their union $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. Then we have $\alpha_{j}^{*}\in\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$.
Proof.
We show that, for every $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, $\alpha_{j}^{\boldsymbol{x}}$ gives the largest revenue of $\boldsymbol{x}+e_{j}$ for ${S}_{j}$ (note that $\boldsymbol{x}+e_{j}=\boldsymbol{x}^{\text{new}}$). Recall that in the main text, we need to increase $\alpha_{j}$ from zero until we find the largest $\alpha_{j}$ such that either of:
  1. 1.
    $g_{k,N,\text{frac}}(\boldsymbol{x}+e_{j})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$,
  2. 2.
    $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})=\overline{\alpha}_{k}$.
If we rewrite the first condition, we are essentially looking for $\alpha_{j}$ such that
$\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})}{u_{k}(\boldsymbol{x}+e_{j})}\right\}$ (43)
Then we see that the revenue that for each $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, seller ${S}_{j}$ can make from buyer $B_{k}$ is
$f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})u_{k}(\boldsymbol{x}_{j}(\boldsymbol{\alpha})),$ (44)
where $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})$ is a non-decreasing function over $\alpha_{j}$ while other terms stays fixed. It indicates that $\alpha_{j}^{\boldsymbol{x}}$ is the largest $\alpha_{j}$ that the seller ${S}_{j}$ could set for buyer $B_{k}$ to purchase $S_{j}$. Therefore, the optimal rate $\alpha_{j}^{*}$ is one of the rates $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. ∎
Appendix E Additional Experimental Details
E.1 Data Valuation Experiments
For each dataset, we randomly sample 200 demonstrations from the validation set to form a representative dataset  [^4] Note: For PIQA we take 200 samples from the training set since the validation set is commonly reserved for testing.. Each data sample in the market is then scored based on its similarity to this representative set.
Model Training: After obtaining purchasing decisions for all data samples, the buyers train their models using the purchased data. In order to conduct a fair comparison across buyers, we sample a set number of data from the buyers’ purchases (shown in Figure˜6). We train each model (i.e., buyer) on these samples separately using LoRA [hu2021loralowrankadaptationlarge] for 3 epochs, with a learning rate of 2e-5 and batch size 32. All models are trained on A6000 GPUs on single GPU settings and take less than 1 hour.
Model Evaluation: For evaluation, we use the test splits of the previously mentioned datasets. In particular, we use 5-shot evaluation on the MathQA test set, and 4-shot evaluation in on the MedQA test. Table˜2 in Appendix˜F shows the demonstrations used for 5-shot and 4-shot evaluation.
E.2 Data Pricing Experiments
Experiment Setups: We simulate two buyer budgets at each time step $t$. The first buyer (high budget) has a budget uniformly randomly generated between $95\%$ and $100\%$ of the total utilities of all $10$ datasets listed in the market. The second buyer (low budget) has a budget uniformly randomly generated between $90\%$ and $95\%$ of the total utilities of all $10$ datasets listed in the market. At each time period, seller’s arriving orders are randomly shuffled. And they prices their own datasets sequentially.
Robustness Check. As discussed in Lemma˜4, the threshold $t^{*}$ when fairshare becomes optimal for the buyer) increases as the discount factor $\delta$ decreases. To run a robustness check, for the high-budget buyer in Figure˜4(a), setting $\delta=0.999,0.99,0.98$ yields $t^{*}=32,38,44$ respectively, which is consistent with our theoretical analysis. In below, we compare the change of buyer’s accumulative utilities over different values of $\delta$.
Figure 7: Buyer’s accumulative utilities over time periods for $\delta=0.980,0.990,0.999$.
Appendix F Additional Experiments and Figures
F.1 Algorithms
Algorithm 1 Determine if buyer $B_{k}$ will purchase dataset $D_{j}$ at price $p_{j}$
1:Inputs: prices $\boldsymbol{p}$, previous optimal $\tilde{\boldsymbol{x}}^{k,N-1}$, feasible set $\mathcal{X}_{k,N-1}$, price $p_{j}$, budget $b_{k}$
2:Output: indicator $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$
3:Initialize $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\leftarrow 0$
[…]
5:for $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$ do
6:  $\boldsymbol{x}^{\text{new}}\leftarrow\boldsymbol{x}+\boldsymbol{e}_{j}$
7:  if $g_{k,N}(\boldsymbol{x}^{\text{new}})>g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$ and $\boldsymbol{x}^{\top}\boldsymbol{p}+p_{j}\leq b_{k}$ then
8:   $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\leftarrow 1$
9:   break
10:  end if
11:end for
12:return $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$
Algorithm 2 Market Dynamic Procedure
1:Inputs: Buyers $\{B_{k}\}_{k=1}^{M}$, Sellers $\{S_{j}\}_{j=1}^{N}$
2:Initialization: Buyers $\{B_{k}\}_{k=1}^{M}$ enter the market
3:for $j=1$ to NN do
4:  Seller $S_{j}$ enters with potential prices $\mathcal{P}_{j,M}$ for dataset $D_{j}$
5:  for all $p_{j}\in\mathcal{P}_{j,M}$ do
6:   for $k=1$ to $M$ do
7:     Buyer $B_{k}$ solves:
$\tilde{\boldsymbol{x}}^{k,j-1}=\arg\max_{\boldsymbol{x}\in\mathcal{X}_{k,j}}g_{k,j}(\boldsymbol{x})$
8:     to decide whether to purchase $D_{j}$ at price $p_{j}$ $\triangleright$ See Eqn. 2
9:   end for
10:   Seller computes net profit $r(p_{j})$ assuming price $p_{j}$ $\triangleright$ See Eqn. 3
11:  end for
12:  Seller selects:
$p_{j}^{*}=\arg\max_{p_{j}\in\mathcal{P}_{j,M}}r(p_{j})$
13:  and sets $p_{j}^{*}$ as the fixed price for $D_{j}$ $\triangleright$ See Eqn. 4
14:end for
F.2 Data Valuation Oracle Experiments
Figure 8: Left, middle-left columns: Correlation analysis between oracle and $\text{Infl}_{\text{IP}}$ valuation. Right, middle-right columns: Performance analysis between between oracle and $\text{Infl}_{\text{IP}}$ valuation.
[…]
To compare the difference between $\text{Infl}_{\text{IP}}$ valuation versus its oracle valuation, we conduct the same experiments described in Section˜4.3. Figure˜8 shows that $\text{Infl}_{\text{IP}}$ and Oracle have decent correlation in their agreement in their valuation of the sellers’ data, which supports findings in previous works [jiao2024context]. We note that in the case where correlation is decent, such as in Figure˜8(b), the final model performance between these methods is close, as seen in Figure˜8(d). In the case where correlation is lower, such as in Figure˜8(a), the final model performance between these methods initially have a gap, but become more similar as the amount of data purchased increases, as seen in Figure˜8(c). This suggests that in practice, even in cases when the agreement between $\text{Infl}_{\text{IP}}$ and Oracle may not be very high, final model performance resulting from these two methods can still be similar.
F.3 Additional Experimental Results
Figure 9: Buyers’ model (top row: Pythia-1b, bottom row: Pythia-410m) performance and costs from their purchased data from math, medical, and physical reasoning data markets. Purchasing decisions were using the constant, random, BM25, $\text{Infl}_{\text{IP}}$ data valuation methods (see Section˜4.3 for details).
Figure 10: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜10(a), 10(e) and 10(i)) and low-budget buyer (Figures˜10(b), 10(f) and 10(j)), and (2) sellers’ average cumulative profits (Figures˜10(c), 10(g) and 10(k)) and number of sellers in the market (Figures˜10(d), 10(h) and 10(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: medqaQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 11: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜11(a), 11(e) and 11(i)) and low-budget buyer (Figures˜11(b), 11(f) and 11(j)), and (2) sellers’ average cumulative profits (Figures˜11(c), 11(g) and 11(k)) and number of sellers in the market (Figures˜11(d), 11(h) and 11(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MathQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 12: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures˜12(a), 12(e) and 12(i)) and low-budget buyer (Figures˜12(b), 12(f) and 12(j)), and (2) sellers’ average cumulative profits (Figures˜12(c), 12(g) and 12(k)) and number of sellers in the market (Figures˜12(d), 12(h) and 12(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: PIQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 13: Number of purchased datasets for the buyer with high budget over time periods ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MedQA, MathQA, and PiQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 14: Number of purchased datasets for the buyer with low budget over time periods ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MedQA, MathQA, and PIQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
F.4 Datasets
Dataset # of Train/Valid/Test Example
MathQA 29837/4475/2985 Question: A train running at the speed of 48 km / hr crosses a pole in 9 seconds . what is the length of the train? a ) 140 , b ) 130 , c ) 120 , d ) 170 , e ) 160
Answer: C
GSM8K 7473/1319 Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May?
Answer: 72
MedQA 10178/1272/1273 Question: A 27-year-old man presents to the emergency room with persistent fever, nausea, and vomiting for the past 3 days. While waiting to be seen, he quickly becomes disoriented and agitated. Upon examination, he has visible signs of difficulty breathing with copious oral secretions and generalized muscle twitching. The patient’s temperature is 104°F (40°C), blood pressure is 90/64 mmHg, pulse is 88/min, and respirations are 18/min with an oxygen saturation of 90% on room air. When the nurse tries to place a nasal cannula, the patient becomes fearful and combative. The patient is sedated and placed on mechanical ventilation. Which of the following is a risk factor for the patient’s most likely diagnosis? a) Contaminated beef b) Epiglottic cyst c) Mosquito bite d) Spelunking
Answer: D
PIQA 16000/2000 Question: How do I ready a guinea pig cage for it’s new occupants? a) Provide the guinea pig with a cage full of a few inches of bedding made of ripped paper strips, you will also need to supply it with a water bottle and a food dish. b) Provide the guinea pig with a cage full of a few inches of bedding made of ripped jeans material, you will also need to supply it with a water bottle and a food dish.
Answer: A
Table 1: Dataset splits and demonstrations from the MathQA, GSM8K, MedQA, and PIQA datasets
Dataset Prompts
MathQA Question: the banker ’ s gain of a certain sum due 3 years hence at 10 % per annum is rs . 36 . what is the present worth ? a ) rs . 400 , b ) rs . 300 , c ) rs . 500 , d ) rs . 350 , e ) none of these
Answer: A
Question: average age of students of an adult school is 40 years . 120 new students whose average age is 32 years joined the school . as a result the average age is decreased by 4 years . find the number of students of the school after joining of the new students . a ) 1200 , b ) 120 , c ) 360 , d ) 240 , e ) none of these
Answer: D
Question: sophia finished 2 / 3 of a book . she calculated that she finished 90 more pages than she has yet to read . how long is her book ? a ) 229 , b ) 270 , c ) 877 , d ) 266 , e ) 281
Answer: B
Question: 120 is what percent of 50 ? na ) 5 % , b ) 240 % , c ) 50% , d ) 2 % , e ) 500
Answer: B
Question: there are 10 girls and 20 boys in a classroom . what is the ratio of girls to boys ? a ) 1 / 2 , b ) 1 / 3 , c ) 1 / 5 , d ) 10 / 30 , e ) 2 / 5
Answer: A
MedQA Question: A mother brings her 3-week-old infant to the pediatrician’s office because she is concerned about his feeding habits. He was born without complications and has not had any medical problems up until this time. However, for the past 4 days, he has been fussy, is regurgitating all of his feeds, and his vomit is yellow in color. On physical exam, the child’s abdomen is minimally distended but no other abnormalities are appreciated. Which of the following embryologic errors could account for this presentation? a) Abnormal migration of ventral pancreatic bud b) Complete failure of proximal duodenum to recanalize c) Abnormal hypertrophy of the pylorus d) Failure of lateral body folds to move ventrally and fuse in the midline
Answer: A
Question: A 53-year-old man comes to the emergency department because of severe right-sided flank pain for 3 hours. The pain is colicky, radiates towards his right groin, and he describes it as 8/10 in intensity. He has vomited once. He has no history of similar episodes in the past. Last year, he was treated with naproxen for swelling and pain of his right toe. He has a history of hypertension. He drinks one to two beers on the weekends. Current medications include amlodipine. He appears uncomfortable. His temperature is 37.10̆0b0C (99.30̆0b0F), pulse is 101/min, and blood pressure is 130/90 mm Hg. Examination shows a soft, nontender abdomen and right costovertebral angle tenderness. An upright x-ray of the abdomen shows no abnormalities. A CT scan of the abdomen and pelvis shows a 7-mm stone in the proximal ureter and grade I hydronephrosis on the right. Which of the following is most likely to be seen on urinalysis? a) Urinary pH: 7.3 b) Urinary pH: 4.7 c) Positive nitrites test d) Largely positive urinary protein
Answer: B
Question: A 48-year-old woman comes to the emergency department because of a photosensitive blistering rash on her hands, forearms, and face for 3 weeks. The lesions are not itchy. She has also noticed that her urine has been dark brown in color recently. Twenty years ago, she was successfully treated for Coats disease of the retina via retinal sclerotherapy. She is currently on hormonal replacement therapy for perimenopausal symptoms. Her aunt and sister have a history of a similar skin lesions. Examination shows multiple fluid-filled blisters and oozing erosions on the forearms, dorsal side of both hands, and forehead. There is hyperpigmented scarring and patches of bald skin along the sides of the blisters. Laboratory studies show a normal serum ferritin concentration. Which of the following is the most appropriate next step in management to induce remission in this patient? a) Pursue liver transplantation b) Begin oral thalidomide therapy c) Begin phlebotomy therapy d) Begin oral hydroxychloroquine therapy
Answer: C
Question: A 23-year-old pregnant woman at 22 weeks gestation presents with burning upon urination. She states it started 1 day ago and has been worsening despite drinking more water and taking cranberry extract. She otherwise feels well and is followed by a doctor for her pregnancy. Her temperature is 97.70̆0b0F (36.50̆0b0C), blood pressure is 122/77 mmHg, pulse is 80/min, respirations are 19/min, and oxygen saturation is 98% on room air. Physical exam is notable for an absence of costovertebral angle tenderness and a gravid uterus. Which of the following is the best treatment for this patient? a) Ampicillin b) Ceftriaxone c) Doxycycline d) Nitrofurantoin
Answer: D
Table 2: Demonstrations included for 5-shot evaluation on the MathQA dataset and for 4-shot evaluation on the MedQA dataset. Demostrations were randomly selected from their respective dataset’s training sets.
Appendix G Limitations and Impact
This paper addresses the critical issue of fairshare pricing in the data market for large language models (LLMs) by proposing a framework and methodologies for fair compensation of datasets from LLM developers to data annotators. Our work directly tackles the ethical and societal challenges in the current data market, where many data annotators are underpaid and receive compensation significantly disconnected from the true economic value their contributions bring to LLMs.
G.1 Limitations
Since our work proposes a novel fairshare framework, there are several lines of future research that can investigate future adjustments to this framework, which lie beyond the scope of our paper. For instance, a large-scale simulation of this market with a wider range of datasets and models is one possibility. In addition, running the simulation with human buyers/sellers is another avenue. Finally, there are several other diverse market dynamics (e.g., incomplete information between buyers/sellers) that can be explored with our proposed framework.
G.2 Impact Statement
From ethical and societal perspectives, our framework prioritizes the welfare of both data annotators and LLM developers. Our methodology ensures that data annotators are fairly compensated for their labor, promoting equity and fairness in the data ecosystem. This contributes to mitigating the exploitation of vulnerable annotators in the data market and aligns the incentives of stakeholders toward a more ethical and sustainable practice. In addition, our framework also benefits LLM developers, by demonstrating that our framework maximizes their utilities and welfare in the long term. Fair compensation encourages ongoing participation of data annotators in the market, ensuring a steady supply of diverse, high-quality datasets essential for LLM development. By addressing existing inequities, our work lays the foundation for a more sustainable, equitable, and mutually beneficial ecosystem for all stakeholders in the LLM data market.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We base on abstract/introduction on the experiments, analysis, and main findings in our paper. Our theoretical and empirical results support the framework we propose.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss our limitations in Appendix˜G
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All assumptions and proofs are discussed in the main paper and in the appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide details on datasets, dataset splits, models, and training procedure/hyperparameters in the paper. In addition, all resources use for our experiments are open-sourced.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our code/data will be openly availiable on GitHub
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We describe all these details in the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The experiments in our paper focus on qualitative or directional insights, not formal statistical claims.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe this in Section˜E.1
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS Code of Ethics and confirms that our research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss this in Appendix˜G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: All data/models used in our research has been obtained from existing open-sourced resources that researchers already use.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We provide citations and acknowledgments for all materials (datasets, models, code) used in our research.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will provide an official open source repo for this paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not conduct any research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not conduct any research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.